OpenStudy (anonymous):

The sum of two consecutive integers is 31. Find the numbers.

OpenStudy (nubeer):

15 and 16 are numbers 15+16 = 31

OpenStudy (aravindg):

take 1st number as n now next number will be n+1 n+(n+1)=31 2n=30 n=15 n+1=16 15 and 16 are the numbers